# Establishing a Foundation for Tetun Text Ad-Hoc Retrieval: Stemming, Indexing, Retrieval, and Ranking

### v2

@@ 8. Labadain-Avaliadór Building
This section provides an overview of the test collection and details the process of constructing Labadain-Avaliadór, a test collection for Tetun. It covers the dataset used, query formulation, document pooling, and relevance judgments.
@@ 8.2. Documents
The document collection consists of 33,550 Tetun documents derived from the Labadain-30k+ dataset (de Jesus and Nunes, 2024b), each enriched with metadata such as title, URL, source, publication date, and content. This dataset was sourced from web crawling and covers a wide range of categories, including news articles, Wikipedia entries, legal and government documents, research papers, technical documents, blogs, forums, and more (de Jesus and Nunes, 2024a). The diversity of its sources and topics makes this dataset particularly suitable for constructing a test collection for Tetun. A sample of the documents, formatted according to TREC guidelines, is shown in Figure 5. The collection comprises 84 MB in size, approximately 12.3 million tokens, and 162,466 unique tokens. A summary of the collection is provided in Table 15 and the length distribution of titles and content is illustrated in Figure 6.
Figure 5. Sample of Document Formatted Following TREC Guidelines: Original (left) and English translation (right).
[…]
Sample of Document Formatted Following TREC Guidelines.
@@
Number of tokens (contents) 11,997,420 2 27,166 357.48 473.99
Figure 6. Length Distribution of Titles and Body Content in the Labadain-30k+ Dataset.
[…]
Length Distribution of Titles and Body Content in the Labadain-30k+ Dataset.
@@ 8.3. Query Formulation
The queries were sourced from the Google Search Console [^8] https://search.google.com/search-console for Timor News [^9] https://www.timornews.tl and from the search logs of users on the Timor News platform. The Google Search Console queries cover the period from November 1, 2021, to January 31, 2024, while the search logs from Timor News span from May 7, 2021, to January 31, 2024. Timor News is an online news agency based in Dili, Timor-Leste, founded in May 2019 and launched its news portal on May 7, 2019. The platform registered an average of 1,400 unique visitors per day and exclusively publishes news in Tetun.
@@
The queries were formulated by five second-year undergraduate volunteers, all native Tetun speakers from Timor-Leste. The group consisted of two students from Environmental Engineering, two from Information Systems, and one from Medicine. They developed the queries according to established guidelines. Each student was initially assigned 250 queries, resulting in a total of 1,250 queries derived from the Google Search Console and Timor News search logs being reviewed. The students analyzed the search logs to understand user search needs, then maintained or modified existing queries or created new ones based on the context provided by the logs.
@@
Before developing the queries, the students participated in a training session that provided practical examples of query formulation. Following this, they participated in three pilot testing sessions. In each session, each student created a query, defined the associated information needs, and described the types of documents that they would consider relevant. These queries were then entered into a search prototype built on top of Apache Solr [^10] https://solr.apache.org using the BM25 ranking model. [^11] https://www.labadain.tl This prototype allowed the students to analyze the documents retrieved for each input query. For each query, each student selected four documents from the retrieved list, ensuring that one document represented each category: irrelevant, marginally relevant, relevant, or highly relevant. These sessions facilitated discussions about the results, highlighted challenges faced, and provided feedback to deepen their understanding of query development.
@@
Figure 7. Sample of Topic Formatted According to TREC Guidelines: Original (left) and English translation (right).
[…]
Sample Topic Formatted According to TREC Guidelines: Original (left) and English Translation (right).
@@
Subsequently, students were tasked to develop short queries following TREC best practices (Sanderson, 2010), ranging from three to five words, specifying information needs and describing the types of documents they expected the system to retrieve. Using the search prototype mentioned earlier, they input their queries and analyzed the retrieved documents. To finalize each query, they ensured that at least five relevant documents could be identified. In total, 61 queries were developed. A sample query (or topic), formatted according to TREC guidelines, is shown in Figure 7. A summary of the queries is provided in Table 16 and their distribution across categories is shown in Figure 8.
@@
Figure 8. Distribution of Queries Over Categories.
[…]
Distribution of Queries Over Categories.
@@ 8.4. Document Pooling
Two probabilistic-based models were used to create the document pool: the $Okapi$ BM25 (known as BM25) and the language model with Dirichlet smoothing. Documents were indexed in separate instances of Solr, each configured to apply one of these ranking models. The document retrieval and pooling process was fully automated and integrated with a relevance assessment interface to streamline the assessment process.
When a query was received, the system retrieved candidate documents ranked by each of the two models, applied balanced interleaving (Radlinski et al., 2008) to combine the results into a pool and presented the top 100 documents to the assessors for relevance judgments. The architecture of the retrieval system used for these assessments is illustrated in Figure 9.
Figure 9. General Architecture of the Retrieval System used for Relevance Judgments.
[…]
General Architecture of the Retrieval System used for Relevance Judgments.
@@ 8.5. Relevance Judgment
Five native Tetun-speaking students who developed the queries conducted relevance judgments for the query-document pairs, categorizing them into four graded levels of topical relevance: irrelevant, marginally relevant, relevant, and highly relevant, as proposed by Sormunen (Sormunen, 2002), following the guidelines outlined in subsection 8.3. A user-friendly web interface was created to streamline the process, allowing assessors to log in with individual accounts to conduct assessments. The interface used for the evaluation is illustrated in Figure 10.
@@
Figure 10. Web Interface Used by Human Assessors for Conducting Relevance Assessments.
[…]
Web Interface Used by Human Assessors for Conducting Relevance Assessments.
Each assessor evaluated the same set of 61 queries, with 100 documents to assess per query, resulting in a total of 6,100 documents being assessed. Assessors were instructed to focus on topical relevance, check document details if the query was too long and not fully displayed on the interface, and disregard the retrieval order when determining the document relevance for the given query. The assessment process was completed within eight hours, and the number of queries assessed by each annotator per hour is illustrated in Figure 11. Documents not included in the judgment list were considered irrelevant.
Figure 11. Hourly Statistics of the Total Number of Queries Evaluated by Each Annotator.
[…]
Hourly Statistics of Queries Evaluated per Annotator.
To assess inter-annotator reliability, we used Cohen’s kappa measure (Cohen, 1960), interpreting the strength of agreement according to the scale provided by Landis and Koch (Landis and Koch, 1977). The overall average score of inter-annotator agreement among the five annotators is 0.4236, indicating moderate agreement, with detailed results presented in Table 17.
@@
An approach to addressing tied scores is to use the tie-breaker strategy, which suggests that using a strong signal to break ties is more effective than a weak one (Wu and Fang, 2013). By applying this tie-breaker method to all instances of tied scores, we obtained the results shown in Table 19 (referring to the 1st round). However, after conducting an in-depth analysis of the tied scores, we observed significant discrepancies in some cases, such as ties between scores of 0 and 2, or 1 and 3. To resolve these inconsistencies, we re-invited three of the five original assessors for a second round of evaluations on the tied documents. During this phase, assessors were presented with the two tied score options from the initial assessment. The reassessment was conducted using Microsoft Excel, with separate tabs for each query and its corresponding documents, and was completed in approximately one hour and 15 minutes.
@@
The final test collection contains an average of 36.76 relevant documents per query, detailed in Table 20, comprising 9.59% highly relevant, 17.86% relevant, 9.31% marginally relevant, and 63.24% irrelevant documents, as shown in the 2nd round column in Table 19. The details of document relevance per topic are illustrated in Figure 12.
Figure 12. Total Number of Relevant and Irrelevant Documents per Topic. Relevant documents consist of marginally relevant, relevant, and highly relevant.
[…]
Total Number of Relevant and Irrelevant Documents per Topic. Relevant documents consist of marginally relevant, relevant, and highly relevant.
8.6. Result and Discussion
Table 19 shows that, after reassessing the tied documents from the first round results using the tie-breaker strategy, the number of highly relevant documents decreased by 144 and relevant documents by 48, while marginally relevant documents increased by 73, and irrelevant documents decreased by 81. This shift suggests that some documents initially classified as highly relevant or relevant were reclassified as marginally relevant, and some marginally relevant documents were reclassified as irrelevant. Additionally, 200 documents from two excluded queries were removed, indicating a change in relevance interpretation after resolving the ties. This reclassification adjusted the distribution of documents across categories, enhancing the overall quality of the test collection.
When analyzing the shifts in relevant documents, annotator 2 exhibited a distinctive pattern, as shown in Figure 11. This annotator demonstrated an unusually high evaluation speed, assessing seven queries (700 documents) in the first hour at an average of 5.14 seconds per query-document pair. This rapid pace is much faster than the average of approximately 8.18 seconds per pair observed among other annotators, who assessed an average of 4.4 queries (440 documents) in the first hour. This rapid pace persisted until the seventh hour, leaving only two queries to be assessed by the eighth hour. Furthermore, Table 17 shows that annotator 2 had only fair agreement with three other annotators (2, 3, and 4), which contributed to increased discrepancies. These observations suggest that the quality of annotations from annotator 2 was lower, probably due to the speed of the evaluation process, which affected the changes in the relevance of the document observed after the second round of evaluations.
@@ 8.6. Result and Discussion
The Labadain-Avaliadór has high relevance density with an average of 36.76 relevant documents per topic—the highest among the collections analyzed. This level of relevance provides a robust foundation for retrieval experiments, supported by a substantial pool of annotated relevance judgments that enable precise and reliable evaluations. Compared to collections with size variations of up to 50,000 documents, such as Sanskrit, and Chichewa, the Labadain-Avaliadór stands out as a well-annotated collection, particularly suited for ad-hoc text retrieval in LRL contexts.
8.7. Conclusions
@@
Approximately 10% of the 6,100 query-document pairs showed disagreement, resulting in tied scores. To resolve these discrepancies, three of the five assessors conducted a second evaluation with only two scoring options for each tied case. This process produced the Labadain-Avaliadór, containing 5,900 qrels, with 9.59% highly relevant documents, 17.86% relevant documents, 9.31% marginally relevant documents, and 63.24% irrelevant documents.
@@ 9. Indexing, Retrieval, and Ranking
This section presents the experiments conducted on ad-hoc text retrieval in Tetun, providing a foundational overview and detailing the retrieval and ranking strategies, as well as the steps involved in text preprocessing, indexing, retrieval, and ranking.
@@ 9.3. Experimental Setting
We explored various retrieval strategies by applying multiple text preprocessing techniques to assess their impact on retrieval effectiveness. The experiment workflow is illustrated in Figure 13. First, we established the baseline by applying the standard preprocessing step, as detailed in subsection 9.2. Next, we tested each of the language-specific text preprocessing techniques, including stopword removal and stemming, to compare their results. Finally, we combined techniques that outperformed the baseline to further evaluate their effectiveness and determine effective strategies for Tetun text ad-hoc retrieval.
Figure 13. Overview of the Experiment Workflow.
[…]
Overview of the Experiment Workflow.
For stemming, although the moderate and heavy variants performed better than the light variant in intrinsic assessment, the difference was minimal. Furthermore, Flores and Moreira (Flores and Moreira, 2016) noted that the most accurate stemmer does not always lead to the most effective retrieval, and therefore we experimented with all stemmer variants.
We used PyTerrier (Macdonald and Tonellotto, 2020), Python API for the Terrier IR platform (Ounis et al., 2005) [^12] http://terrier.org/ for indexing, retrieval, and ranking, with the default settings maintained for each model in the experiments. The same text preprocessing techniques were applied to both queries and documents, and experiments were conducted using document titles and content to evaluate retrieval effectiveness for each approach.
@@ 9.4. Document Indexing
We indexed document titles and content independently, with the index created from standard preprocessing steps serving as the baseline. Additionally, we indexed individual results from applying language-specific preprocessing steps, stopword removal, and each of the stemmer variants. To assess the effects of these preprocessing methods on index size, index compression factors (ICF) (Frakes and Fox, 2003) were calculated for each preprocessing step. A summary of the index compression results is provided in Table 22.
The results show that all preprocessing methods generally reduce index size compared to the baseline. Removing hyphens yielded the highest compression factor for title indexing, reducing the index size by up to 30.76% compared to the baseline. For content indexing, the heavy stemmer provided the most significant compression, reducing the index size by up to 12.18%.
@@
As expected, the moderate stemmer variant compressed the index more than the light variant, and the heavy variant achieved the most compression for both title and content indexing. Interestingly, removing apostrophes and accents also reduced index size, with accent removal achieving more than a 7% reduction for both titles and content, suggesting a significant presence of identical words with and without accents in the documents.

### v6

@@ 8. Labadain-Avaliadór Building
This section provides an overview of the test collection and details the process of constructing Labadain-Avaliadór (avaliadór, a Tetun word meaning “evaluator”), a Tetun test collection for evaluation. It covers the dataset used, query formulation, document pooling, and relevance judgments.
@@ 8.2. Documents
The document collection comprises 33,550 Tetun documents sourced from the Labadain-30k+ dataset (de Jesus and Nunes, 2024b), each enriched with metadata such as title, URL, source, publication date, and content. This dataset was collected from web crawling and covers a wide range of categories, including news articles, Wikipedia entries, legal and government documents, research papers, technical documents, blogs, forums, and more (de Jesus and Nunes, 2024a). The diversity of its sources and topics makes this dataset particularly suitable for constructing a test collection for Tetun. A sample of the documents, formatted according to TREC guidelines, is shown in Figure 5. The collection is 84 MB in size, with approximately 12.3 million tokens and 162,466 unique tokens. A summary of the collection is provided in Table 15, and the length distribution of titles and content is illustrated in Figure 6.
Figure 5. Sample of Document Formatted Following TREC Guidelines: Original (left) and English translation (right).:
@@
Number of tokens (content) 11,997,420 2 27,166 357.48 473.99
Figure 6. Length Distribution of Titles and Body Content in the Labadain-30k+ Dataset. In the right-hand figure, the x-axis of the document length distribution is limited to 1,000 tokens (words and numbers) for improved visualization.:
@@ 8.3. Query Formulation
Queries were collected from two sources: Google Search Console [^8] https://search.google.com/search-console for Timor News and the user search logs from the Timor News platform. [^9] https://www.timornews.tl The Google Search Console queries cover the period from November 1, 2021, to January 31, 2024, while the search logs from Timor News span from May 7, 2021, to January 31, 2024. Timor News is an online news agency based in Dili, Timor-Leste, founded in May 2019 and launched its news portal on May 7, 2019. The platform registered an average of 1,400 unique visitors per day and exclusively publishes news in Tetun.
@@
The collected queries were compiled and distributed among five second-year undergraduate volunteers, all native Tetun speakers from Timor-Leste. The group comprised two students from Environmental Engineering, two from Information Systems, and one from Medicine. These students were tasked with developing queries following the established guidelines. Initially, each student was assigned 250 queries, resulting in a total of 1,250 queries analyzed. To understand user information needs, the students reviewed the provided search logs and either retained, modified, or formulated new queries based on the contextual information interpreted from the logs (see examples in Table 16).
Table 16. Examples of original query logs and their reformulations. Words highlighted in green background indicate newly added terms, while word in orange background indicates orthography correction.
Original query log Reformulated query
Problema lixu (Waste problem) Problema lixu iha Dili (Waste problem in Dili)
Soe bebe (Baby abandonment) Kazu soe bebé (Case of baby abandonment)
Konsumu tabaku (Tobacco consumption) Dadus konsumu tabaku (Tobacco consumption data)
Before beginning query development, the students attended a training session that provided practical examples of query formulation. Following this, three pilot testing sessions were conducted, during which each student created a query, defined the associated information need, and described the types of documents they would consider relevant.
@@
The queries were then entered into a search prototype [^10] https://www.labadain.tl built on top of Apache Solr [^11] https://solr.apache.org using the BM25 ranking model. This prototype allowed the students to analyze the documents retrieved for each input query. For each query, each student selected four documents from the top 50 retrieved list, ensuring that one document represented each category: non-relevant, marginally relevant, relevant, or highly relevant. These sessions facilitated discussions about the results, highlighted challenges faced, and provided feedback to deepen their understanding of query development.
@@
Figure 7. Sample of Topic Formatted According to TREC Guidelines: Original (left) and English translation (right).:
@@
Subsequently, students were tasked to develop short queries following TREC best practices (Sanderson, 2010), ranging from three to five words, specifying information needs and describing the types of documents they expected the system to retrieve. Using the search prototype mentioned earlier, they input their queries and analyzed the retrieved documents. To finalize each query, they ensured that at least five relevant documents were identified. In total, 61 queries were developed. A sample query (or topic), formatted according to TREC guidelines, is shown in Figure 7. A summary of the queries is provided in Table 17, and their distribution across categories is shown in Figure 8. For categorization, we adapted the query topic categorization frameworks of Beitzel et al. (2004) and Rohatgi et al. (2021).
@@
Figure 8. Distribution of Queries Over Categories.:
@@ 8.4. Document Pooling
Since Query-document relevance judgments are carried out by human assessors, it is not feasible to evaluate every document in a large collection. To address this challenge, Spärck-Jones and van Rijsbergen (1975) introduced the pooling technique, in which a small subset containing a sufficiently representative sample of relevant documents is selected from the larger collection and provided to human assessors for relevance judgments (Sanderson, 2010).
Given the limited availability of retrieval models and techniques for LRLs, and to maximize the retrieval of relevant documents for constructing a robust test collection, we created the document pool using two retrieval models: BM25 and a language model (LM) with Dirichlet smoothing. BM25 is widely recognized for its effectiveness in ad-hoc retrieval (Robertson and Zaragoza, 2009), while the LM with Dirichlet smoothing has been shown to perform particularly well on short queries (Zhai and Lafferty, 2001). To balance the contributions of these models, we applied the balanced interleaving technique (Radlinski et al., 2008) to merge their results into a pool, which was then presented to assessors for evaluation.
Documents were indexed in separate instances of Solr, each configured with either the BM25 or the Dirichlet LM ranking models. The document retrieval and pooling process was fully automated and integrated with a relevance assessment interface to streamline the workflow. When a query was received, the system retrieved candidate documents ranked by each of the two models, merged the results into a pool, and presented the top 100 documents to the assessors for relevance judgments. The architecture of the retrieval system used for these assessments is shown in Figure 9.
Figure 9. General Architecture of the Retrieval System used for Relevance Judgments. The document highlighted in red indicates a duplicate that was excluded from the final list.:
@@ 8.5. Relevance Judgment
Five native Tetun-speaking students who developed the queries conducted relevance judgments for the query-document pairs, categorizing them into four graded levels of topical relevance: non-relevant, marginally relevant, relevant, and highly relevant, as proposed by Sormunen (2002), following the guidelines outlined in Subsection 8.3. A user-friendly web interface was created to streamline the process, allowing assessors to log in with individual accounts to conduct assessments. The interface used for the evaluation is illustrated in Figure 10.
Figure 10. Web Interface Used by Human Assessors for Conducting Relevance Assessments.:
@@
Each assessor evaluated the same set of 61 queries, with 100 documents to assess per query, resulting in a total of 6,100 documents being assessed. Assessors were instructed to focus on topical relevance, check document details if the query was too long and not fully displayed on the interface, and disregard the retrieval order when determining the relevance of the document for the given query. The assessment process was completed within eight hours, and the number of queries assessed by each annotator per hour is illustrated in Figure 11. Documents not included in the judgment list were considered non-relevant.
Figure 11. Hourly Statistics of the Total Number of Queries Evaluated by Each Annotator.:
To assess inter-annotator reliability, we used Cohen’s kappa measure (Cohen, 1960), interpreting the strength of agreement according to the scale provided by Landis and Koch (1977). The overall average score of inter-annotator agreement among the five annotators is 0.4236, indicating moderate agreement, with detailed results presented in Table 18.
@@
An approach to addressing tied scores is to use the tie-breaker strategy, which suggests that using a strong signal to break ties is more effective than a weak one (Wu and Fang, 2013). By applying this tie-breaker method to all instances of tied scores, we obtained the results shown in Table 20 (referring to the 1st round). However, after conducting an in-depth analysis of the tied scores, we observed significant discrepancies in some cases, such as ties between scores of 0 and 2 or 1 and 3. To resolve these inconsistencies, we re-invited three of the five original assessors for a second round of evaluations on the tied documents. During this phase, assessors were presented with the two tied score options from the initial assessment. The reassessment was conducted using Microsoft Excel, with separate tabs for each query and its corresponding documents, and was completed in approximately one hour and 15 minutes.
@@
The final test collection, called Labadain-Avaliadór (de Jesus and Nunes, 2025a), contains an average of 36.76 relevant documents per query, detailed in Table 21—comprising 9.59% highly relevant, 17.86% relevant, 9.31% marginally relevant, and 63.24% non-relevant documents—as shown in the 2nd round column in Table 20. The distribution of document relevance per query is illustrated in Figure 12.
Figure 12. Total Number of Relevant and Non-relevant Documents per Query. Relevant documents consist of marginally relevant, relevant, and highly relevant.:
8.6. Results and Discussion
Table 20 shows that, after reassessing the tied documents from the first round results using the tie-breaker strategy, the number of highly relevant documents decreased by 144 and relevant documents by 48, while marginally relevant documents increased by 73, and non-relevant documents decreased by 81. This shift suggests that some documents initially classified as highly relevant or relevant were reclassified as marginally relevant, and some marginally relevant documents were reclassified as non-relevant. Additionally, 200 documents from two excluded queries were removed, indicating a change in relevance interpretation after resolving the ties. This reclassification adjusted the distribution of documents across categories, enhancing the overall quality of the test collection.
When analyzing the changes in document relevance, annotator 2 exhibited a distinct pattern, as illustrated in Figure 11. This annotator judged seven queries (700 documents) in the first hour at an average rate of 5.14 seconds per query-document pair, significantly faster than the average of approximately 8.18 seconds per pair observed among other annotators, who assessed an average of 4.4 queries (440 documents) in the first hour. This fast pace persisted until the seventh hour, leaving only two queries for the eighth hour. Additionally, Table 18 shows that annotator 2 had only fair agreement with three other annotators (2, 3, and 4), which contributed to increased discrepancies. These observations suggest that the quality of annotations from annotator 2 was lower, probably due to the speed of the evaluation process, which affected the changes in the relevance of the document observed after the second round of evaluations.
@@ 8.6. Results and Discussion
The Labadain-Avaliadór has high relevance density with an average of 36.76 relevant documents per topic, the highest among the collections analyzed. This level of relevance provides a robust foundation for retrieval experiments, supported by a substantial pool of annotated relevance judgments that enable precise and reliable evaluations. Compared to collections with size variations of up to 50,000 documents, such as Sanskrit and Chichewa, the Labadain-Avaliadór stands out as a well-annotated collection, particularly suited for ad-hoc text retrieval in LRL contexts.
8.7. Conclusion
@@
Approximately 10% of the 6,100 query-document pairs showed disagreement, resulting in tied scores. To resolve these discrepancies, three of the five assessors conducted a second evaluation with only two scoring options for each tied case. This process produced the Labadain-Avaliadór, containing 5,900 qrels, with 9.59% highly relevant documents, 17.86% relevant documents, 9.31% marginally relevant documents, and 63.24% non-relevant documents.
@@ 9. Indexing, Retrieval, and Ranking
This section presents the experiments conducted on Tetun ad-hoc text retrieval, providing a detailed description of the retrieval and ranking strategies and the steps involved in text preprocessing, indexing, retrieval, and ranking.
@@ 9.3. Experimental Setting
We explored various retrieval strategies by applying multiple text preprocessing techniques to assess their impact on retrieval effectiveness. The experiment workflow is illustrated in Figure 13. First, we established the baseline by applying the standard preprocessing step, as detailed in Subsection 9.2. Next, we tested each of the language-specific text preprocessing techniques, including stopword removal and stemming, to compare their results. Finally, we combined techniques that outperformed the baseline to further evaluate their effectiveness and determine effective strategies for Tetun ad-hoc text retrieval.
Figure 13. Overview of the Experiment Workflow.:
For stemming, although the moderate and heavy variants performed better than the light variant in intrinsic assessment, the difference was minimal. Furthermore, Flores and Moreira (2016) noted that the most accurate stemmer does not always lead to the most effective retrieval, and therefore, we experimented with all stemmer variants. We used PyTerrier (Macdonald and Tonellotto, 2020), a Python API for the Terrier IR platform (Ounis et al., 2005) [^12] http://terrier.org/ for indexing, retrieval, and ranking, with the default settings maintained for each model. The same text preprocessing techniques were applied to both queries and documents, and experiments were conducted using document titles and content to evaluate retrieval effectiveness for each approach.
@@ 9.4. Document Indexing
Document titles and content were indexed independently, with the index created from standard preprocessing steps serving as the baseline. Additional indexes were generated for individual results obtained from applying language-specific preprocessing steps, stopword removal, and each of the stemmer variants. To evaluate the effects of these preprocessing methods on index size, index compression factors (ICF) (Frakes and Fox, 2003) were calculated for each preprocessing configuration. A summary of the index compression results is presented in Table 23.
The results show that all preprocessing methods generally reduce the index size compared to the baseline. Removing hyphens yielded the highest compression factor for title indexing, reducing the index size by up to 30.76% compared to the baseline. For content indexing, the heavy stemmer provided the highest compression, reducing the index size by up to 12.18%.
@@
As expected, the moderate stemmer variant compressed the index more efficiently than the light variant, while the heavy variant achieved the highest compression for both title and content indexing. Interestingly, removing apostrophes and accents also contributed to index size reduction, with accent removal achieving over a 7% reduction for both titles and content. This suggests a substantial presence of identical words with and without accents in the documents.